OpenStudy (anonymous):

Kathleen and Elissa were repairing Kathleen's club house. Kathleen carried a board from her family's garage to the club house and leaned the board against a side wall of the club house. The board leaned against the wall with the bottom of the board six feet from the wall and the top of the board touching the wall at eight feet above ground. How long, in feet, was the board Kathleen carried?

OpenStudy (anonymous):

8^2+6^2=C^2 C=10